In an attempt to determine the factors that affect the amount of energy used, 200 households were analysed. In each, the number of occupants and the amount of electricity used (watt-hours) were measured. The estimated linear regression equation describing the relationship between the Amount of energy used (y) and the Number of occupants (x) is given by y^ = 458.9 + 64x. Forecast the amount of energy used (in watt-hours) in a household with 4 occupants (x = 4):

a.

458.9

b.

714.9

c.

64

d.

2091.6
Answer #1

Solution:

The estimated linear regression equation is

  y^ = 458.9 + 64x

Put x = 4 in this equation.

   y^ = 458.9 + (64 * 4)

= 458.9 + 256

= 714.9

Answer is option b. 714.9